### v1

# There is no universal separable Banach algebra

Abstract.
We show that no separable Banach algebra is universal for homomorphic embeddings of all separable Banach algebras, whether embeddings are merely bounded or required to be contractive. The commutative version also fails: no separable commutative Banach algebra is universal for embeddings of all separable commutative algebras.
The proofs follow the same pattern but use different linearisation spaces. Given a bounded bilinear form $\beta$, we associate a separable test algebra $A(\beta)$ whose multiplication records $\beta$. Any homomorphic embedding of $A(\beta)$ forces the linearisation of $\beta$ to factor through a fixed separable space: in the commutative case through $R\widehat{\otimes}_{\pi,s}R$ (where $R$ is the Jacobson radical of the target algebra), and in the general case through $B\widehat{\otimes}_{\pi}B$. Choosing $\beta$ so that the corresponding operator fails to factor through that space, using the theorem of Johnson–Szankowski, yields a contradiction. The non‑commutative argument avoids radicals entirely and uses only the ordinary projective tensor product.
Key words and phrases:
Banach algebra, Jacobson radical, symmetric projective tensor product, universal Banach algebra, Johnson–Szankowski theorem
2020 Mathematics Subject Classification:
46H10, 46H15, 46M05
RVO: 67985840.
1. Introduction
A natural question in the theory of Banach algebras concerns the existence of universal objects for certain classes thereof: does there exist a separable Banach algebra that contains (in a suitable sense) all separable Banach algebras from the given class? The answer may depend critically on both the algebraic structure considered and the notion of ‘containment’ employed.
For commutative C*-algebras, the situation is well understood. Since every compact metrisable space is a quotient of the Cantor set $\Delta:=2^{\mathbb{N}}$, the algebra $C(\Delta)$ of continuous functions on the Cantor set is universal for separable commutative C*-algebras: every such algebra embeds isometrically as a subalgebra of $C(\Delta)$. In contrast, there is no separable C*-algebra that is universal for all separable C*-algebras (commutative or not), as can be deduced from [5, Proposition 2.6].
For general Banach algebras without additional structure, it is folklore that there exists no separable commutative Banach algebra $B$ such that every separable commutative Banach algebra admits an isometric embedding into $B$. A straightforward proof appears in [6]: if $p$ and $q$ are commuting projections in a Banach algebra, then $\|p-q\|\geqslant 1$, so any set of commuting projections is discrete; since projections in Banach algebras can have arbitrarily large norms, a commutative Banach algebra containing uncountably many commuting projections must be non-separable.
However, the question of universality for bounded (or contractive) homomorphic embeddings, i.e., embeddings that are merely required to have closed range, appears to have been left open in the literature. This is the gap we address in the present work.
Our main result, Theorem 1.1, shows that there is no separable (commutative) Banach algebra $B$ that is universal for embeddings of separable (commutative) Banach algebras, both in the category of contractive homomorphisms and in the category of (arbitrary) bounded homomorphisms. The proofs combine constructions using projective tensor products (symmetric in the commutative case, ordinary in the general case) with the deep Johnson–Szankowski theorem [4, Theorem B] that the class of compact operators does not factor through any separable Banach space. By constructing test algebras whose radical structure forces any homomorphic embedding to induce a factorisation of a carefully chosen non-factorable compact operator, we obtain the desired contradiction.
We work over the field of complex numbers however the proofs are valid for real Banach algebras too. A homomorphism between Banach algebras means a bounded algebra homomorphism. An embedding is an injective homomorphism that is a topological isomorphism onto its (closed) range. Sometimes we additionally require homomorphisms to be contractive; in the other we do not, however we always stress this out explicitly.
Theorem 1.1.
There are no separable universal objects for the categories of separable Banach algebras, in either the commutative or the general (not-necessarily commutative) setting, for contractive or merely bounded homomorphic embeddings. More precisely,
  1. (A)
    There is no separable commutative Banach algebra $B$ such that every separable commutative Banach algebra $A$ admits a contractive homomorphic embedding $A\hookrightarrow B$.
  2. (B)
    There is no separable commutative Banach algebra $B$ such that every separable commutative Banach algebra $A$ admits a (bounded) homomorphic embedding $A\hookrightarrow B$.
  3. (C)
    There is no separable Banach algebra $B$ such that every separable Banach algebra $A$ admits a contractive homomorphic embedding $A\hookrightarrow B$.
  4. (D)
    There is no separable Banach algebra $B$ such that every separable Banach algebra $A$ admits a (bounded) homomorphic embedding $A\hookrightarrow B$.
We present a single proof that treats the commutative and general cases separately but in either case covers both morphism kinds at once. Contractivity plays no rôle beyond boundedness.
2. Preliminaries
We refer to [7, Chapter 2] for all facts related to the projective tensor product of Banach spaces and to [1, Chapter 2] for Banach-algebraic aspects thereof. For a Banach space $X$ and an integer $n$, the ($n$-fold) projective tensor product $X^{\widehat{\otimes}_{\pi}n}$ (or simply $X\widehat{\otimes}_{\pi}X$ for $n=2$) is defined as the completion of the algebraic tensor product $X^{\otimes n}$ under the projective norm.
Theorem 2.1 (Universality of the projective tensor product).
Let $X,Y$ be Banach spaces and $n\in\mathbb{N}$. For every continuous $n$-linear map $M\colon X^{n}\to Y$ there exists a unique bounded operator $L_{M}\in\mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi}n},\,Y\big)$ such that
$M=L_{M}\circ\otimes^{n}_{\pi}.$
Moreover, the correspondence $M\mapsto L_{M}$ is an isometric linear isomorphism
$\mathcal{L}(X^{n};Y)\ \cong\ \mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi}n},\,Y\big),$
so that $\|L_{M}\|=\|M\|$. Diagrammatically,
[…]
Lemma 2.2 (Functoriality of $\widehat{\otimes}_{\pi}$).
Let $u:X\to X_{1}$ and $v:Y\to Y_{1}$ be bounded linear maps between Banach spaces. There is a unique bounded operator
$u\otimes_{\pi}v:\ X\widehat{\otimes}_{\pi}Y\ \longrightarrow\ X_{1}\widehat{\otimes}_{\pi}Y_{1}$
such that $(u\otimes_{\pi}v)(x\otimes y)=u(x)\otimes v(y)$ for all $x\in X$, $y\in Y$, and $\|u\otimes_{\pi}v\|\leqslant\|u\|\,\|v\|$. In particular, if $B$ is a Banach algebra with multiplication $\mu_{B}:B\times B\to B$ and linearisation $\widetilde{\mu}_{B}:B\widehat{\otimes}_{\pi}B\to B$, then for any $u_{1},u_{2}:X\to B$ we have
$\widetilde{\mu}_{B}\circ(u_{1}\otimes_{\pi}u_{2})\;=\;L_{\mu_{B}\circ(u_{1}\times u_{2})},$
the (unique) linearisation of the bilinear map $(x,x^{\prime})\mapsto\mu_{B}(u_{1}(x),u_{2}(x^{\prime}))$. An analogous statement holds for the symmetric projective tensor product.
Proof.
The well-defininedness of the operator $u_{1}\otimes_{\pi}u_{2}$ is standard consequence of the universality of the projective tensor product and the norm estimate follows from reasonability of the projective crossnorm. For the identity, evaluate both sides on elementary tensors:
$\bigl(\widetilde{\mu}_{B}\circ(u_{1}\otimes_{\pi}u_{2})\bigr)(x\otimes x^{\prime})=\widetilde{\mu}_{B}\big(u_{1}(x)\otimes u_{2}(x^{\prime})\big)=\mu_{B}\big(u_{1}(x),u_{2}(x^{\prime})\big).$
By uniqueness in Theorem 2.1, this determines the linearisation. ∎
Lemmas 2.2 and 2.3 provide the basic functorial calculus for linearisations that we shall use repeatedly. The key point is that if two bilinear maps $M$ and $N$ are related by precomposition with linear maps, their linearisations are related by the induced tensor product maps.
Lemma 2.3 (Compatibility of linearisation).
Let $M:X_{1}\times Y_{1}\to Z$ be a bounded bilinear map and let $u:X\to X_{1}$, $v:Y\to Y_{1}$ be bounded linear maps. Define $N:=M\circ(u\times v):X\times Y\to Z$. Then
$\widetilde{N}\;=\;\widetilde{M}\circ(u\otimes_{\pi}v).$
In particular, if $M=N$ as bilinear maps on the same domain, then $\widetilde{M}=\widetilde{N}$.
Lemma 2.4.
Let $u:X\to Z$ be bounded. Then there is a unique bounded operator
$u\otimes_{\mathrm{sym}}u:\ X\widehat{\otimes}_{\pi,s}X\ \longrightarrow\ Z\widehat{\otimes}_{\pi,s}Z$
with $(u\otimes_{\mathrm{sym}}u)(x\odot x^{\prime})=u(x)\odot u(x^{\prime})$ on symmetric tensors, and
$\|\,u\otimes_{\mathrm{sym}}u\,\|\ \leqslant\ \|u\|^{2}.$
Proof.
Let $q_{X}:X\widehat{\otimes}_{\pi}X\to X\widehat{\otimes}_{\pi,s}X$ and $q_{Z}:Z\widehat{\otimes}_{\pi}Z\to Z\widehat{\otimes}_{\pi,s}Z$ be the canonical quotient maps (of norm $1$). Set $T:=q_{Z}\circ(u\otimes_{\pi}u)$; then $\|T\|\leqslant\|u\|^{2}$ by Lemma 2.2. By the universal property of $q_{X}$ there exists a unique $U:X\widehat{\otimes}_{\pi,s}X\to Z\widehat{\otimes}_{\pi,s}Z$ with $U\circ q_{X}=T$; this is $u\otimes_{\mathrm{sym}}u$. For $z\in X\widehat{\otimes}_{\pi,s}X$ and $\varepsilon>0$ choose $w$ with $q_{X}(w)=z$ and $\|w\|\leqslant\|z\|+\varepsilon$. Then
$\|Uz\|=\|Tw\|\leqslant\|T\|\,\|w\|\leqslant\|u\|^{2}(\|z\|+\varepsilon);$
letting $\varepsilon\downarrow 0$ proves the claim. ∎
For a Banach space $X$ and an integer $n$, the ($n$-fold) symmetric projective tensor product $X^{\widehat{\otimes}_{\pi,s}n}$ (or simply $X\widehat{\otimes}_{\!\pi,s}X$ for $n=2$) is defined as the completion of the algebraic symmetric tensor product $X^{\odot n}$ under the projective norm inherited from $X^{\widehat{\otimes}_{\pi}n}$. Here $X^{\odot n}$ denotes the quotient of the algebraic tensor product $X^{\otimes n}$ by the subspace generated by all elements of the form $x_{1}\otimes\cdots\otimes x_{n}-x_{\sigma(1)}\otimes\cdots\otimes x_{\sigma(n)}$, with $\sigma$ ranging over the symmetric group $S_{n}$. It provides a natural linearisation space for continuous symmetric $n$-linear maps.
Theorem 2.5 (Universality of the symmetric projective tensor product).
Let $X,Y$ be Banach spaces and $n\in\mathbb{N}$. For every continuous symmetric $n$-linear map $M\colon X^{n}\to Y$ there exists a unique bounded operator $L_{M}\in\mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi,s}n},\,Y\big)$ such that
$M=L_{M}\circ\otimes^{n}_{s},\qquad\text{equivalently}\quad M=(L_{M}\circ\sigma^{n}_{X})\circ\otimes^{n}_{\pi}.$
Here $\otimes^{n}_{\pi}\colon X^{n}\to X^{\widehat{\otimes}_{\pi}n}$ is the canonical $n$-linear map, $\sigma^{n}_{X}\colon X^{\widehat{\otimes}_{\pi}n}\to X^{\widehat{\otimes}_{\pi,s}n}$ is the canonical quotient onto the symmetric projective tensor product, and $\otimes^{n}_{s}:=\sigma^{n}_{X}\circ\otimes^{n}_{\pi}$. Diagrammatically,
[…]
Moreover, the correspondence $M\mapsto L_{M}$ is an isometric linear isomorphism
$\mathcal{L}_{s}(X^{n};Y)\ \cong\ \mathcal{L}\!\big(X^{\widehat{\otimes}_{\pi,s}n},\,Y\big),$
so that $\|L_{M}\|=\|M\|$. In particular, $(X^{\widehat{\otimes}_{\pi,s}n})^{*}$ is isometrically the space of continuous $n$-homogeneous polynomials on $X$.
Proof.
This is standard; see Hájek–Johanis [2, Theorem 13]. ∎
Remark 2.6.
For any bounded bilinear $\beta:X\times X\to Y$ we write $\widetilde{\beta}:X\widehat{\otimes}_{\pi}X\to Y$ for its (unsymmetrised) projective linearisation. If $\beta$ is symmetric, we also write $\widetilde{\beta}^{\,\mathrm{sym}}:X\widehat{\otimes}_{\pi,s}X\to Y$ for the symmetric linearisation. In the commutative parts of the paper we use $\widetilde{\beta}^{\,\mathrm{sym}}$ consistently.
If $R$ is a commutative Banach algebra, its multiplication $\mu_{R}:R\times R\to R$ is symmetric and linearises to $\widetilde{\mu}_{R}^{\mathrm{sym}}:R\widehat{\otimes}_{\!\pi,s}R\to R$.
Lemma 2.7 (Image of the radical under homomorphisms).
Let $\phi:A\to B$ be a bounded homomorphism of Banach algebras. Then $\phi\bigl(\operatorname{rad}(A)\bigr)\subseteq\operatorname{rad}\bigl(\overline{\phi(A)}\bigr)$.
Proof.
This is standard: the Jacobson radical is functorial for surjective maps and, in general, one has $\phi(\operatorname{rad}A)\subseteq\operatorname{rad}(\overline{\phi(A)})$; see, e.g., [1, §3.3]. ∎
Lemma 2.8.
Let $A$ be a Banach algebra and let $I\triangleleft A$ be a (two-sided) ideal with $I^{n}=0$ for some $n\in\mathbb{N}$. Then $I\subseteq\operatorname{rad}(A)$. Moreover, if $A/I$ is semisimple (equivalently $\operatorname{rad}(A/I)=0$), then $\operatorname{rad}(A)\subseteq I$.
Proof.
If $a\in I$ then $(1-a)(1+a+\dots+a^{n-1})=1$, so 1a1-a is invertible in $A$. Hence $a$ lies in the Jacobson radical. The second assertion follows from the standard functoriality of the radical under quotient maps, see [1, §3.3]. ∎
We now construct our test algebras. Given a bounded bilinear map $\beta:X\times X\to Y$, we build a three-level algebra $A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y$ where the scalar level $\mathbb{C}$ acts as an identity, the middle level $X$ acts on itself via $\beta$ to produce elements in the top level $Y$, and the top level $Y$ is multiplicatively inert (squares to zero). The multiplication thus ‘records’ the bilinear form $\beta$. Any homomorphism embedding $A(\beta)$ into another algebra will be forced to preserve this structure, leading to factorisation constraints that we exploit via the Johnson–Szankowski theorem.
Proposition 2.9.
Let $X,Y$ be Banach spaces, and let $\beta:X\times X\to Y$ be a bounded bilinear map. Define a multiplication on the $\ell_{1}$-direct sum
$A(\beta):=\mathbb{C}\oplus_{1}X\oplus_{1}Y$
by
$(\alpha,x,y)\cdot(\alpha^{\prime},x^{\prime},y^{\prime})=\bigl(\alpha\alpha^{\prime},\ \alpha x^{\prime}+\alpha^{\prime}x,\ \alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr).$
Then $A(\beta)$ is a unital Banach algebra (after an equivalent renorming if desired). It is commutative whenever $\beta$ is symmetric, and in all cases
$\operatorname{rad}\bigl(A(\beta)\bigr)\;=\;X\oplus Y.$
If $\beta$ is symmetric, it linearises to a unique bounded operator $\widetilde{\beta}:X\widehat{\otimes}_{\pi,\mathrm{s}}X\to Y$.
Proof.
We equip $A(\beta)$ with the norm $\|(\alpha,x,y)\|=|\alpha|+\|x\|+\|y\|$.
First, we verify that the multiplication is bounded. Take $a=(\alpha,x,y)$ and $b=(\alpha^{\prime},x^{\prime},y^{\prime})$ in $A(\beta)$. Then
$\|ab\|$ $=|\alpha\alpha^{\prime}|+\|\alpha x^{\prime}+\alpha^{\prime}x\|+\|\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\|$
$\leqslant|\alpha\alpha^{\prime}|+|\alpha|\,\|x^{\prime}\|+|\alpha^{\prime}|\,\|x\|+|\alpha|\,\|y^{\prime}\|+|\alpha^{\prime}|\,\|y\|+\|\beta\|\,\|x\|\,\|x^{\prime}\|.$
Now observe that $|\alpha\alpha^{\prime}|\leqslant|\alpha|\,\|b\|$ and $|\alpha|\,\|x^{\prime}\|\leqslant\|a\|\,\|b\|$; similarly for the other terms. A routine calculation shows that
$\|ab\|\leqslant(5+\|\beta\|)\,\|a\|\,\|b\|.$
If desired, one may renorm $A(\beta)$ by scaling the given norm by the constant factor $5+\|\beta\|$ to make the multiplication contractive.
The element $e=(1,0,0)$ is clearly a two-sided identity for the multiplication.
We now establish associativity. Let $a=(\alpha,x,y)$, $b=(\alpha^{\prime},x^{\prime},y^{\prime})$, and $c=(\alpha^{\prime\prime},x^{\prime\prime},y^{\prime\prime})$. We have
$ab=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr),$
whence
$(ab)c=\Bigl((\alpha\alpha^{\prime})\alpha^{\prime\prime},\,(\alpha\alpha^{\prime})x^{\prime\prime}+\alpha^{\prime\prime}(\alpha x^{\prime}+\alpha^{\prime}x),\,S\Bigr),$
where
$S$ $=(\alpha\alpha^{\prime})y^{\prime\prime}+\alpha^{\prime\prime}(\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime}))+\beta(\alpha x^{\prime}+\alpha^{\prime}x,x^{\prime\prime})$
$=\alpha\alpha^{\prime}y^{\prime\prime}+\alpha\alpha^{\prime\prime}y^{\prime}+\alpha^{\prime}\alpha^{\prime\prime}y+\alpha^{\prime\prime}\beta(x,x^{\prime})+\alpha\beta(x^{\prime},x^{\prime\prime})+\alpha^{\prime}\beta(x,x^{\prime\prime}),$
by the bilinearity of $\beta$. On the other hand,
$bc=\bigl(\alpha^{\prime}\alpha^{\prime\prime},\,\alpha^{\prime}x^{\prime\prime}+\alpha^{\prime\prime}x^{\prime},\,\alpha^{\prime}y^{\prime\prime}+\alpha^{\prime\prime}y^{\prime}+\beta(x^{\prime},x^{\prime\prime})\bigr),$
so that
$a(bc)=\Bigl(\alpha(\alpha^{\prime}\alpha^{\prime\prime}),\,\alpha(\alpha^{\prime}x^{\prime\prime}+\alpha^{\prime\prime}x^{\prime})+(\alpha^{\prime}\alpha^{\prime\prime})x,\,T\Bigr),$
where
$T$ $=\alpha(\alpha^{\prime}y^{\prime\prime}+\alpha^{\prime\prime}y^{\prime}+\beta(x^{\prime},x^{\prime\prime}))+(\alpha^{\prime}\alpha^{\prime\prime})y+\beta(x,\alpha^{\prime}x^{\prime\prime}+\alpha^{\prime\prime}x^{\prime})$
$=\alpha\alpha^{\prime}y^{\prime\prime}+\alpha\alpha^{\prime\prime}y^{\prime}+\alpha\beta(x^{\prime},x^{\prime\prime})+\alpha^{\prime}\alpha^{\prime\prime}y+\alpha^{\prime}\beta(x,x^{\prime\prime})+\alpha^{\prime\prime}\beta(x,x^{\prime}).$
Comparing coordinates, we see that the first two coordinates of $(ab)c$ and $a(bc)$ coincide immediately, while the third coordinates $S$ and $T$ are manifestly equal (they consist of the same six terms). Thus the multiplication is associative. (We remark that only the bilinearity of $\beta$ is used here; symmetry plays no rôle.)
If $\beta$ is symmetric, commutativity follows. Indeed,
$ab=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr)=\bigl(\alpha^{\prime}\alpha,\,\alpha^{\prime}x+\alpha x^{\prime},\,\alpha^{\prime}y+\alpha y^{\prime}+\beta(x^{\prime},x)\bigr)=ba.$
Next, we determine the radical of $A(\beta)$. Set $J=X\oplus Y=\{(0,x,y):x\in X,\,y\in Y\}$. It is straightforward to check that $J$ is an ideal of $A(\beta)$. Moreover, $J$ is nilpotent: for $(0,x,y)\in J$, we have
$(0,x,y)^{2}=(0,0,\beta(x,x))\quad\text{and}\quad(0,x,y)^{3}=0.$
Since $J^{3}=\{0\}$, Lemma 2.8 gives $J\subseteq\operatorname{rad}(A(\beta))$. The quotient $A(\beta)/J\cong\mathbb{C}$ is semi-simple, so again by Lemma 2.8 we have $\operatorname{rad}(A(\beta))\subseteq J$. Hence
$\operatorname{rad}\bigl(A(\beta)\bigr)=X\oplus Y\$}.
This argument uses only that $J$ is nilpotent, so it holds whether or not $\beta$ is symmetric.
Finally, we observe that the symmetric bilinear map $\beta:X\times X\to Y$ linearises, by Theorem 2.5, to a unique bounded linear operator $\widetilde{\beta}:X\widehat{\otimes}_{\pi,\mathrm{s}}X\to Y$ satisfying $\widetilde{\beta}(x\otimes_{\mathrm{sym}}x^{\prime})=\beta(x,x^{\prime})$ for all $x,x^{\prime}\in X$, with $\|\widetilde{\beta}\|=\|\beta\|$. ∎
We now show that when $A(\beta)$ embeds into a commutative Banach algebra $B$, the image must respect the radical structure. Since $\operatorname{rad}(A(\beta))=X\oplus Y$ is nilpotent and homomorphisms preserve radicals in the commutative setting, the embedding forces $X$ and $Y$ to land in $\operatorname{rad}(B)$. The multiplication structure in $A(\beta)$ then translates into a factorisation identity through the symmetric tensor product of the radical.
Lemma 2.10.
Let $B$ be a commutative Banach algebra with radical $R=\operatorname{rad}(B)$, and let $X$ and $Y$ be Banach spaces. Consider the algebra
$A(\beta):=\mathbb{C}\oplus_{1}X\oplus_{1}Y,\qquad(\alpha,x,y)\cdot(\alpha^{\prime},x^{\prime},y^{\prime})=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr),$
where $\beta:X\times X\to Y$ is a bounded symmetric bilinear map, and let $\widetilde{\beta}^{\,\mathrm{sym}}:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to Y$ be its unique bounded linearisation through the symmetric projective tensor product. Suppose that $\phi:A(\beta)\to B$ is a bounded algebra homomorphism, and write
$i:=\phi|_{X}:X\longrightarrow R,\qquad j:=\phi|_{Y}:Y\longrightarrow R.$
Then
(2.1) $j\circ\widetilde{\beta}^{\,\mathrm{sym}}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(\,i\otimes_{\mathrm{sym}}i\,)\quad:\quad X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\longrightarrow R,$
where $\mu_{R}:R\times R\to R$ denotes multiplication in the radical $R$, $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R\to R$ is its symmetric linearisation, and $i\otimes_{\mathrm{sym}}i:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R$ is the bounded map induced by $i$ on symmetric tensor products. Moreover, if $\phi$ is a homomorphic embedding, then both $i$ and $j$ are injective, and $j$ is bounded below. Hence $j$ (and similarly $i$) is injective with closed range and admits a bounded inverse on its range.
Proof.
By Proposition 2.9, $J:=X\oplus Y$ is an ideal of $A(\beta)$ with $J^{3}=0$, hence every $z\in J$ is nilpotent. Therefore $\phi(z)$ is nilpotent in $B$ and lies in $R:=\operatorname{rad}(B)$. In particular the restrictions
$i:=\phi|_{X}:X\to R,\qquad j:=\phi|_{Y}:Y\to R$
are well-defined and bounded.
For each $x\in X$, write $\mathbf{x}:=(0,x,0)\in A(\beta)$, and for each $y\in Y$ write $\mathbf{y}:=(0,0,y)\in A(\beta)$. The multiplication in $A(\beta)$ gives
$\mathbf{x}\cdot\mathbf{x}^{\prime}\ =\ (0,0,\beta(x,x^{\prime}))\qquad(x,x^{\prime}\in X).$
Applying the homomorphism $\phi$ to this identity, and noting that $\phi(\mathbf{x})=i(x)$, $\phi(\mathbf{x}^{\prime})=i(x^{\prime})$, and $\phi(0,0,y)=j(y)$ all lie in $R$, we obtain
(2.2) $j\bigl(\beta(x,x^{\prime})\bigr)\ =\ \phi(\mathbf{x}\cdot\mathbf{x}^{\prime})\ =\ \phi(\mathbf{x})\,\phi(\mathbf{x}^{\prime})\ =\ \mu_{R}\bigl(i(x),i(x^{\prime})\bigr)$
for all $x,x^{\prime}\in X$.
As $\beta$ is a bounded symmetric bilinear map, we may apply Theorem 2.5 with $Z=Y$ to obtain a unique bounded operator $\widetilde{\beta}:X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\to Y$ satisfying $\widetilde{\beta}(x\odot_{\mathrm{s}}x^{\prime})=\beta(x,x^{\prime})$ for all $x,x^{\prime}\in X$. Similarly, since $R$ is commutative, the multiplication $\mu_{R}:R\times R\to R$ is a bounded symmetric bilinear map. Applying Theorem 2.5 with $Z=R$, we obtain a unique bounded operator $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ R\to R$ such that $\widetilde{\mu}_{R}^{\,\mathrm{sym}}(u\odot_{\mathrm{s}}v)=\mu_{R}(u,v)$ for all $u,v\in R$.
Furthermore, by Lemma 2.7, the homomorphism $\phi:A(\beta)\to B$ maps $\operatorname{rad}(A(\beta))$ into the radical $R=\operatorname{rad}(B)$, so its restriction $i=\phi|_{X}:X\to R$ is a bounded linear map. The functoriality of the symmetric projective tensor product then yields a canonical bounded operator
$i\otimes_{\mathrm{sym}}i\colon X\widehat{\otimes}_{\pi,\mathrm{s}}X\longrightarrow R\widehat{\otimes}_{\pi,\mathrm{s}}R,$
defined on elementary symmetric tensors by
$(i\otimes_{\mathrm{sym}}i)(x\odot_{\mathrm{s}}x^{\prime})=i(x)\odot_{\mathrm{s}}i(x^{\prime}),\qquad x,x^{\prime}\in X,$
and extended to the completion by continuity. The map $i\mapsto i\otimes_{\mathrm{sym}}i$ is continuous and one has the norm estimate $\|\,i\otimes_{\mathrm{sym}}i\,\|\leqslant\|i\|^{2}$.
Consider the bounded operators
$T_{1}:=j\circ\widetilde{\beta}:\ X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\ \to\ R,\qquad T_{2}:=\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i):\ X\ \widehat{\otimes}_{\!\pi,\mathrm{s}}\ X\ \to\ R.$
For all $x,x^{\prime}\in X$, we have by construction and by (2.2):
$T_{1}(x\odot_{\mathrm{s}}x^{\prime})\$$=\ j\bigl(\beta(x,x^{\prime})\bigr)\ =\ \mu_{R}\bigl(i(x),i(x^{\prime})\bigr)$
$=\ \widetilde{\mu}_{R}^{\,\mathrm{sym}}\bigl(i(x)\odot_{\mathrm{s}}i(x^{\prime})\bigr)\ =\ T_{2}(x\odot_{\mathrm{s}}x^{\prime}).$
[…]
Suppose now that $\phi$ is an embedding, so that $\phi$ is injective and admits a bounded inverse $\psi:\phi(A(\beta))\to A(\beta)$ on its range. Then, for each $y\in Y$, we have
$\|y\|\ =\ \bigl\|\psi\bigl(\phi(0,0,y)\bigr)\bigr\|\ \leqslant\ \|\psi\|\,\bigl\|\phi(0,0,y)\bigr\|\ =\ \|\psi\|\,\|j(y)\|.$
Consequently, $\|j(y)\|\geqslant\|\psi\|^{-1}\|y\|$ for all $y\in Y$, which shows that $j$ is bounded below. In particular, $j$ is injective. An entirely analogous argument applied to $x\in X$ establishes that $i=\phi|_{X}$ is injective (and likewise bounded below). ∎
The proof works the same for the unsymmetrised version.
Lemma 2.11.
Let $B$ be a (possibly noncommutative) Banach algebra, and let $X,Y$ be Banach spaces. For a bounded bilinear map $\beta:X\times X\to Y$ let $\widetilde{\beta}:X\widehat{\otimes}_{\!\pi}X\to Y$ denote its linearisation. If $\phi:A(\beta)\to B$ is a bounded homomorphism and we set
$i:=\phi|_{X}:X\to B,\qquad j:=\phi|_{Y}:Y\to B,$
then
$j\circ\widetilde{\beta}\;=\;\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i)\quad:\quad X\widehat{\otimes}_{\!\pi}X\longrightarrow B,$
where $\mu_{B}:B\times B\to B$ is multiplication and $\widetilde{\mu}_{B}:B\widehat{\otimes}_{\!\pi}B\to B$ its linearisation. If $\phi$ is an embedding, then $j$ is injective and bounded below.
Proof.
Let $M:=\mu_{B}\circ(i\times i)$ and $N:=j\circ\beta$ as bounded bilinear maps on $X\times X$. From $(0,x,0)(0,x^{\prime},0)=(0,0,\beta(x,x^{\prime}))$ in $A(\beta)$ and multiplicativity of $\phi$ we get $M(x,x^{\prime})=N(x,x^{\prime})$ for all $x,x^{\prime}$. By Lemma 2.3 we obtain $\widetilde{M}=\widetilde{N}$. Using Lemma 2.2, $\widetilde{M}=\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i)$, while $\widetilde{N}=j\circ\widetilde{\beta}$, yielding the identity. If $\phi$ is an embedding with inverse $\psi$ on its range, then $\|y\|=\|\psi(\phi(0,0,y))\|\leqslant\|\psi\|\,\|j(y)\|$, so $j$ is bounded below. ∎
Remark 2.12.
Note that Lemma 2.11 does not use any radical structure: $i$ and $j$ need not land in a distinguished ideal of $B$. All factorisations proceed through the fixed separable space
$G:=B\widehat{\otimes}_{\pi}B$
via the canonical map $i\otimes_{\pi}i:E\to G$ and the linearised multiplication $\widetilde{\mu}_{B}:G\to B$.
To apply the Johnson–Szankowski theorem, we must carefully choose the bilinear form $\beta$ so that its linearisation $\widetilde{\beta}$ does not factor through a given separable space $G$. The construction proceeds in two steps: first, we build a surjective bilinear form $\beta_{U}$ with a bounded right inverse (at the linearised level), and then compose with a compact operator $S:U\to Y$ that fails to factor through $G$. The surjectivity ensures that non-factorisation of the composition implies non-factorisation of $S$, completing the circle.
Lemma 2.13.
Let $G$ be a separable Banach space, and let $U$ and $Y$ be Banach spaces. Suppose that $S:U\to Y$ is a compact operator which does not factor through $G$. Write $X:=U\oplus_{1}\mathbb{C}$, and define a symmetric bilinear map $\beta_{U}:X\times X\to U$ by
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big)=\lambda u^{\prime}+\lambda^{\prime}u$
for $(u,\lambda),(u^{\prime},\lambda^{\prime})\in X$. Let $E:=X\widehat{\otimes}_{\pi,\mathrm{s}}X$, and let $\widetilde{\beta}_{U}:E\to U$ be the linearisation of $\beta_{U}$. Then $\widetilde{\beta}_{U}$ is surjective and admits a bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$. Consequently, writing $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, we have:
  • $T_{0}$ is compact, and
  • $T_{0}$ does not factor through $G$ (for otherwise $S=T_{0}\circ\sigma$ would factor through $G$, contradicting our hypothesis).
Proof.
For each $u\in U$, observe that
$\widetilde{\beta}_{U}\big((u,0)\otimes_{\mathrm{sym}}(0,1)\big)=\beta_{U}\big((u,0),(0,1)\big)=u.$
Thus $\widetilde{\beta}_{U}$ is surjective, and the map $\sigma:U\to E$ defined by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$ satisfies $\widetilde{\beta}_{U}\circ\sigma=\operatorname{id}_{U}$.
To verify that $\sigma$ is bounded, for elementary symmetric tensors in any symmetric projective tensor product, one has
$\big\|x\otimes_{\mathrm{sym}}y\big\|_{\pi,\mathrm{s}}\leqslant\|x\|\,\|y\|,$
because $X\widehat{\otimes}_{\!\pi,s}X$ is a quotient of the projective tensor product that is a reasonable crossnorm. Consequently,
$\|\sigma(u)\|_{\pi,\mathrm{s}}\ \leqslant\ \|(u,0)\|\,\|(0,1)\|\ =\ \|u\|\cdot 1\ =\ \|u\|,$
so that $\sigma$ is indeed bounded (in fact, a contraction).
Since $S$ is compact and $\widetilde{\beta}_{U}$ is bounded, their composition $T_{0}=S\circ\widetilde{\beta}_{U}$ is compact. Suppose, towards a contradiction, that $T_{0}$ factors through $G$, say $T_{0}=Q\circ P$ with $P:E\to G$ and $Q:G\to Y$ both bounded. Then
$S\ =\ T_{0}\circ\sigma\ =\ Q\circ(P\circ\sigma),$
which exhibits $S$ as factoring through $G$, contrary to hypothesis. Thus $T_{0}$ does not factor through $G$. ∎
An analogous fact holds true in the non-symmetric setting.
Lemma 2.14.
Let $G$ be a separable Banach space, and let $U,Y$ be Banach spaces. Suppose $S:U\to Y$ is a compact operator which does not factor through $G$. Set $X:=U\oplus_{1}\mathbb{C}$ and define a bilinear map
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda\,u^{\prime}\qquad\big((u,\lambda),(u^{\prime},\lambda^{\prime})\in X\big).$
Let $E:=X\widehat{\otimes}_{\pi}X$, and let $\widetilde{\beta}_{U}:E\to U$ be its linearisation. Then $\widetilde{\beta}_{U}$ is surjective and admits a bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(0,1)\otimes(u,0)$. Consequently, writing $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, we have:
  • $T_{0}$ is compact; and
  • $T_{0}$ does not factor through $G$ (otherwise $S=T_{0}\circ\sigma$ would factor through $G$).
Proof.
For each $u\in U$,
$\widetilde{\beta}_{U}\big((0,1)\otimes(u,0)\big)=\beta_{U}\big((0,1),(u,0)\big)=u,$
so $\widetilde{\beta}_{U}$ is surjective and $\sigma(u):=(0,1)\otimes(u,0)$ is a right inverse. Since $\|(0,1)\otimes(u,0)\|_{\pi}\leqslant\|(0,1)\|\,\|(u,0)\|=\|u\|$, the map $\sigma$ is bounded. Compactness of $T_{0}=S\circ\widetilde{\beta}_{U}$ follows because $S$ is compact and $\widetilde{\beta}_{U}$ is bounded. If $T_{0}$ factored through $G$ as $Q\circ P$ then $S=T_{0}\circ\sigma=Q\circ(P\circ\sigma)$ would factor through $G$, contradicting the hypothesis. ∎
Proof of Theorem 1.1.
First, we present a single proof that establishes both assertions (A) and (B) simultaneously. Let $B$ be an arbitrary separable commutative Banach algebra, and write $R:=\operatorname{rad}(B)$ for its Jacobson radical. Since $B$ is separable, so too is $R$. Set
$G:=R\widehat{\otimes}_{\pi,\mathrm{s}}R,$
the symmetric projective tensor product of $R$ with itself. Then $G$ is a separable Banach space.
By the Johnson–Szankowski theorem [4, Theorems 2.5–2.6], there exist separable Banach spaces $U$ and $Y$ and a compact operator $S:U\to Y$ which does not factor through the separable space GG. [^1] By a consequence of Johnson–Szankowski [4, Thm. 2.5] (see also [3] for background), for every separable Banach space $G$ there exist separable Banach spaces $U$ and $Y$ and a compact operator $S:U\to Y$ such that $S$ does not belong to the operator ideal of maps factoring through $G$ (i.e., $S$ does not factor through $G$).
We now construct a separable commutative test algebra that cannot embed into $B$. Write $X:=U\oplus_{1}\mathbb{C}$ (the $\ell_{1}$-direct sum), and define a symmetric bilinear map $\beta_{U}:X\times X\to U$ by
$\beta_{U}\big((u,\lambda),(u^{\prime},\lambda^{\prime})\big):=\lambda u^{\prime}+\lambda^{\prime}u$
for all $(u,\lambda),(u^{\prime},\lambda^{\prime})\in X$. Let $E:=X\widehat{\otimes}_{\pi,\mathrm{s}}X$, and let $\widetilde{\beta}^{\,\mathrm{sym}}_{U}:E\to U$ be the unique bounded linearisation of $\beta_{U}$. By Lemma 2.13, the map $\widetilde{\beta}_{U}$ is surjective with bounded right inverse $\sigma:U\to E$ given by $\sigma(u):=(u,0)\otimes_{\mathrm{sym}}(0,1)$, and the compact operator
$T_{0}:=S\circ\widetilde{\beta}^{\,\mathrm{sym}}_{U}:E\longrightarrow Y$
does not factor through $G$ (for if it did, then $S=T_{0}\circ\sigma$ would factor through $G$, contradicting our choice of $S$). Now define a symmetric bilinear map $\beta:X\times X\to Y$ by composition:
$\beta:=S\circ\beta_{U}.$
Then the linearisation of $\beta$ is precisely $\widetilde{\beta}^{\,\mathrm{sym}}=S\circ\widetilde{\beta}^{\,\mathrm{sym}}_{U}=T_{0}$. By Proposition 2.9, the $\ell_{1}$-direct sum
$A:=A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y,$
equipped with the multiplication
$(\alpha,x,y)\cdot(\alpha^{\prime},x^{\prime},y^{\prime}):=\bigl(\alpha\alpha^{\prime},\,\alpha x^{\prime}+\alpha^{\prime}x,\,\alpha y^{\prime}+\alpha^{\prime}y+\beta(x,x^{\prime})\bigr),$
is a commutative, unital Banach algebra (possibly after an equivalent renorming) with radical $\operatorname{rad}(A)=X\oplus Y$. Since $U$, $Y$, and hence $X$ are all separable, the algebra $A$ is separable.
Suppose, towards a contradiction, that there exists an embedding $\phi:A\hookrightarrow B$. (In case (A), we require $\phi$ to be contractive; in case (B), we merely require $\phi$ to be bounded. In either case, $\phi$ is an injective homomorphism that is a topological isomorphism onto its closed range in $B$.) Write
$i:=\phi|_{X}:X\longrightarrow R\qquad\text{and}\qquad j:=\phi|_{Y}:Y\longrightarrow R$
for the restrictions of $\phi$ to $X$ and $Y$, respectively. By Lemma 2.10, we have the factorisation identity
(2.3) $j\circ\widetilde{\beta}^{\,\mathrm{sym}}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i)\quad:\quad E=X\widehat{\otimes}_{\pi,\mathrm{s}}X\longrightarrow R,$
where $\widetilde{\mu}_{R}^{\,\mathrm{sym}}:R\widehat{\otimes}_{\pi,\mathrm{s}}R\to R$ is the symmetric linearisation of multiplication in $R$. Moreover, since $\phi$ is an embedding, Lemma 2.10 also tells us that $j:Y\to R$ is bounded below (and in particular injective). Thus $j$ is a topological isomorphism of $Y$ onto the closed subspace $\operatorname{ran}(j)\subseteq R$, and the inverse map $(j|_{\operatorname{ran}j})^{-1}:\operatorname{ran}(j)\to Y$ is bounded.
Using (2.3) and $\widetilde{\beta}^{\,\mathrm{sym}}=T_{0}$ we have
$j\circ T_{0}\;=\;\widetilde{\mu}_{R}^{\,\mathrm{sym}}\circ(i\otimes_{\mathrm{sym}}i)\ :\ E\longrightarrow R,$
so $j\circ T_{0}$ factors through $G=R\widehat{\otimes}_{\pi,s}R$ (via $E\xrightarrow{i\otimes_{\mathrm{sym}}i}G\xrightarrow{\ \widetilde{\mu}_{R}^{\,\mathrm{sym}}\ }R$). By Lemma 2.13 we have a bounded right inverse $\sigma:U\to E$ of $\widetilde{\beta}^{\,\mathrm{sym}}_{U}$ with $S=T_{0}\circ\sigma$, hence
$j\circ S\;=\;(j\circ T_{0})\circ\sigma$
also factors through $G$. Since $\phi$ is an embedding, $j:Y\to R$ is injective and bounded below, so the inverse $(j|_{\operatorname{ran}j})^{-1}:\operatorname{ran}j\to Y$ is bounded. Regard $j\circ S$ as an operator with codomain $\operatorname{ran}j$ (a Banach space with the inherited norm). Post-composing by $(j|_{\operatorname{ran}j})^{-1}$ yields
$S\;=\;(j|_{\operatorname{ran}j})^{-1}\circ(j\circ S),$
and factorisations are stable under bounded post-composition. Therefore $S$ also factors through $G$, contradicting the choice of $S$. This completes the proof of (A) and (B).
The argument for parts (C) and (D) follows the same overall strategy but with two crucial differences: we work with the unsymmetrised projective tensor product $B\widehat{\otimes}_{\pi}B$, and we avoid any reliance on radical structure. The test algebra construction uses a non-symmetric bilinear form, and the factorisation proceeds directly through multiplication in $B$.
Fix a separable Banach algebra $B$ and set $G:=B\widehat{\otimes}_{\pi}B$, which is separable. Again, by the Johnson–Szankowski theorem, there exist separable spaces $U,Y$ and a compact operator $S:U\to Y$ which does not factor through $G$. Form $X:=U\oplus_{1}\mathbb{C}$ and $\beta_{U}$ as in Lemma 2.14; let $E:=X\widehat{\otimes}_{\pi}X$ and $T_{0}:=S\circ\widetilde{\beta}_{U}:E\to Y$, which does not factor through $G$. Define $\beta:=S\circ\beta_{U}:X\times X\to Y$ and let $A:=A(\beta)=\mathbb{C}\oplus_{1}X\oplus_{1}Y$, a separable Banach algebra by Proposition 2.9. Assuming $B$ is universal, there is an embedding $\phi:A\hookrightarrow B$. Write $i:=\phi|_{X}$ and $j:=\phi|_{Y}$. By Lemma 2.11,
$j\circ T_{0}\;=\;\widetilde{\mu}_{B}\circ(i\otimes_{\pi}i).$
Set $P:=i\otimes_{\pi}i:E\to G$ and $Q:=\widetilde{\mu}_{B}:G\to B$; then
$j\circ T_{0}=Q\circ P,$
so $j\circ T_{0}$ factors through $G=B\widehat{\otimes}_{\pi}B$. By Lemma 2.14 we have $j\circ S=(j\circ T_{0})\circ\sigma=Q\circ(P\circ\sigma)$, hence $j\circ S$ also factors through $G$. Since $\phi$ is an embedding, $j$ is injective and bounded below; viewing $j\circ S$ with codomain $\operatorname{ran}j$ and post-composing with $(j|_{\operatorname{ran}j})^{-1}$ shows that $S$ factors through $G$, a contradiction. ∎
Remark 2.15.
The compactness of the operator $S:U\to Y$ is not essential for the argument. It is used only so that we may appeal directly to the Johnson–Szankowski theorem in its classical form, which states that no separable Banach space is universal for the factorisation of compact operators. The proof requires merely that $S$ be a bounded operator between separable Banach spaces which does not belong to the operator ideal of maps factoring through $G:=R\widehat{\otimes}_{\pi,\mathrm{s}}R.$ In particular, one could take $S=\mathrm{id}_{U}$ for a separable Banach space $U$ which does not embed complementably into $G$; this again relies on the earlier result of Johnson and Szankowski [3], asserting that no separable Banach space contains complemented copies of all separable Banach spaces. We retain the compact case for definiteness.
Acknowledgements
Funding received from NCN Sonata-Bis 13 (2023/50/E/ST1/00067) is acknowledged with thanks.